OpenStudy (anonymous):

(sqrt 7x) times ((sq rt x)-7sq rt.7) is the answer x times sq rt 7 -49 times sq rt. x?

OpenStudy (anonymous):

\[\sqrt{7x}(\sqrt{x}-7\sqrt{7})=\sqrt{7}-7(7)\sqrt{x}=\sqrt{7}-49\sqrt{x}\]

OpenStudy (anonymous):

not x times sq rt. 7?

OpenStudy (anonymous):

don't I have x squared =x

OpenStudy (anonymous):

no because\[\sqrt{7x}\sqrt{x}=\sqrt{7}\sqrt{x}\sqrt{x}=x \sqrt{7}\]

OpenStudy (anonymous):

that's what I thought but you don't have x in front of sq rt 7 in your answer

OpenStudy (jiteshmeghwal9):

\[\sqrt{7x} \times (\sqrt{x}-7\sqrt{7})\]\[(\sqrt{7x} \times \sqrt{x})(\sqrt{7x} \times 7\sqrt{7})\]

OpenStudy (anonymous):

is my answer not correct?

OpenStudy (jiteshmeghwal9):

ohh ! i don't see that @Jonask has already helped u :)

OpenStudy (anonymous):

why not x times sq rt.7 as part of answer